An optically active mixture has an ee value of 60% ee. What is percentage of the major enantiomerin this mixture?A. 75%B. 80%C. 60%D. 40%
Answer #1

Opti (B ) is correct option

% ce = (+) A% - (-) A % Suppos (+) A% is n sthan (-)A% is (100-1) Hence 60 = n- (100-1) = n-100th 160 = 2n =) N=80 (+) A% = 8